# From Lengthy to Lucid: A Systematic Literature Review on NLP Techniques for Taming Long Sentences

### v1

@@ 3.2.1. Extractive Datasets
The first dataset for extractive sentence compression is extracted from the Ziff-Davis corpus (Knight and Marcu, 2000). Ziff-Davis has a relatively small corpus of newspaper articles announcing computer products. A small portion of the corpus has about 1,000 sentence-compression pairs (Knight and Marcu, 2002b). This dataset has been mostly used for training and evaluating early works on sentence compression. In addition,  Vandeghinste and Tjong Kim Sang (2004) build a parallel corpus for subtitle generation that contains transcriptions from several broadcasts. The British National Corpus (BNC) and the English Broadcast News Corpus are two other small human-annotated extractive sentence compression datasets that are typically used for evaluation purposes (Clarke and Lapata, 2008). The first one comprises 82 newspaper articles (1,433 sentences) from the BNC and the American News Text corpus, while the second one comprises 50 stories (1,370 sentences) from the HUB-4 1996 English Broadcast News corpus.
@@
Summarization datasets have also been used for the evaluation of sentence compression. For example, a small extractive dataset of 280 examples is constructed in (Ghalandari et al., 2022) based on Newsroom. Newsroom (Grusky et al., 2018) is a large dataset for training and evaluating summarization systems. It contains 1.3 million articles and summaries written by authors and editors in the newsrooms of 38 major publications. The summaries are obtained from search and social metadata between 1998 and 2017 and use a variety of summarization strategies combining extraction and abstraction. Ghalandari et al. (2022) extract the first three sentences from each article with a length between 15 and 60 tokens. They consider cases where the tokenized summary was contained in a tokenized sentence. They further filter these cases based on grammaticality and informativeness.
@@
Google sentence compression (Filippova and Altun, 2013) is the largest dataset for extractive sentence compression. It is automatically constructed, consisting of 250K first sentences of a news article along with their extractive compression derived from information in the corresponding article’s headline [^1] https://github.com/google-research-datasets/sentence-compression.
3.2.2. Abstractive Datasets
@@
Abstractive datasets contain sentence-compression pairs, where the compression is not necessarily a direct sub-sequence of the original long sentence. While the majority of related literature has focused on the news domain, there also exists work in other domains, such as subtitles or Wikipedia documents. For example, Yamangil and Nelken (2008) exploit Wikipedia revision history to enrich compression data with more than 380,000 compression pairs.
@@
A small abstractive dataset is constructed in (Rush et al., 2015) based on the data of Task 1 and 2 of the 2004 Document Understanding Conferences (DUC) [^2] https://duc.nist.gov/duc2004/. It contains 500 news articles from the New York Times and Associated Press Wire services, each paired with four human summaries. A repository containing these data is available [^3] https://github.com/UsmanNiazi/DUC-2004-Dataset. For the purpose of sentence summarization, the first sentence of each article is used, while summaries are clipped at 75 characters. This leads to summaries with an average of 10 words.
@@
A larger abstractive text compression dataset (Toutanova et al., 2016) is introduced from Microsoft Research (MSR). It consists of 26,000 long sentences accompanied by their respective compressions. Each long sentence is paired with multiple target compressions (1 to 5) rated by quality judgments regarding sentence coherence and meaning preservation. Besides the newswire domain, the MSR dataset also includes journals as well as business and technical reports. One of the largest datasets for abstractive sentence compression is Gigaword (Napoles et al., 2012). Gigaword is used in (Rush et al., 2015) to create a dataset by pairing the first sentence of each article with its headline. Gigaword is a large-scale dataset that is used for summarization and headline generation. It contains approximately 9.5 million articles from different news sources. After rule-based filtering of some spurious pairs, about 4 million examples remained.
@@
Abstractive datasets for sentence compression are either used only for evaluation (DUC-2004) while larger datasets such as Gigaword (Napoles et al., 2012) and MSR (Toutanova et al., 2016) are also used for training and evaluation.
3.3. Measures
@@
3.3.1. Compression Rate
One of the most widely used metric for evaluating sentence compression methods is Compression Rate (Knight and Marcu, 2000; McDonald, 2006; Galley and McKeown, 2007; Galanis and Androutsopoulos, 2010; Filippova et al., 2015; Zhou and Rush, 2019; Zhao et al., 2018; Ghalandari et al., 2022). It is defined as the number of words in the compressed text divided by the number of tokens in the uncompressed text as follows:
(1) Compression Rate=# words in compressed sentence# words in original $\text{Compression Rate}=\frac{\text{\# words in compressed sentence}}{\text{\#words in original sentence}}$
@@ 3.3.2. ROUGE score
The family of ROUGE metrics (Lin, 2004) is one of the most popular measures for text summarization. Since text summarization and sentence compression share many similarities in terms of shortening the initial text into a more concise form, ROUGE is also widely used for sentence compression. ROUGE metrics are based on the n-gram overlaps or the longest common subsequence (LCS) between the target and the prediction and can be used to evaluate the quality of the generated text. ROUGE consists of several metrics such as ROUGE-N (ROUGE-1, ROUGE-2, etc.), ROUGE-L, and ROUGE-S. In sentence compression, ROUGE metrics are computed between the predicted and ground truth compressions. More specifically, ROUGE-1 and ROUGE-2 are used for measuring the overlap of unigrams and bigrams between the predicted and ground truth compression, respectively. ROUGE-L is used for measuring the LCS between the generated and the target compression.
@@
Even though ROUGE is an established evaluation metric for sentence compression, Zhang et al. (2021) highlight some of its limitations. For example, high ROUGE score might be misleading producing factual errors. Li et al. (2018) reveal that existing encoder-decoder models (Zhou et al., 2017) that achieve significant performance according to established metrics like ROUGE score might not generate faithful compressions. Another important factor in this context is the informativeness of the generated compressions. Zheng et al. (2020) work towards this direction by proposing an entity-controllable compression model that ensures the presence of the most relevant entities of the input sentence in the compressions. This method works by first extracting all the entities from the input sentence and ranking the most important ones with an entity selector module. Finally, instead of conventional left-to-right decoding, the model decodes the text around the selected entities, ensuring that the resulting compression is both concise and informative.
3.3.3. Other measures
@@
Other measures include simple token-overlapping (Ghalandari et al., 2022), grammatical relations (Clarke and Lapata, 2006b; Filippova and Altun, 2013) or even human evaluation (Filippova and Altun, 2013). In specific, Clarke and Lapata (2006b) demonstrate that measuring the grammatical relations between the input and the target is a measure that correlates well with human judgments. The F1 score between the functional structures of the sentence and its compression, obtained via Lexical-Functional Grammar parsing, was proposed in (Riezler et al., 2003) and used in (Filippova and Altun, 2013). In addition, simple F1 score between the tokens of the predicted and ground truth compression are computed in (Ghalandari et al., 2022). In (Filippova and Altun, 2013), humans were shown pairs of sentences and their compression and were asked to rate in a scale from 1 to 5 the readability (i.e., grammaticallity, fluency) and the informativeness (i.e., importance, representatives) of the compression. Three ratings were collected for every compression.
3.4. Comparative Evaluation
@@
Table 3. Results of sentence compression methods on different extractive datasets. We report ROUGE-1 (R-1), ROUGE-2 (R-2) and ROUGE-L (R-L) scores, F-measure, relations F-measure (Relations F1) and compression rate (CR). All scores are reported in %. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
Ziff-Davis (Knight and Marcu, 2000) - - - 68.80 - 70.19
(Unno et al., 2006) - - - 78.58 - 58.14
(Clarke and Lapata, 2008) - - - - 46.77 48.67
(Cohn and Lapata, 2009) - - - - 56.55 67.45
Google (Filippova et al., 2015) - - - 82.0 - 38.0
(Zhao et al., 2018) - - - 85.1 - 39.0
(Kamigaito and Okumura, 2020) 79.3 71.4 79.1 85.5 - -
(Park et al., 2021) - - - 86.2 - 40.9
BNC (Clarke and Lapata, 2008) - - - - 50.1 68.5
(Thadani and McKeown, 2013) - - - 71.91 - -
(Thadani, 2014) - - - 74.00 - -
(Ghalandari et al., 2022) 79.49 62.32 78.63 76.50 - 76.0
Broadcast (Clarke and Lapata, 2008) - - - 40.8 63.7
(Ghalandari et al., 2022) 82.22 66.01 81.78 78.7 - 78.0
(Thadani and McKeown, 2013) - - - 77.82 - -
(Thadani, 2014) - - - 80.13 - -
@@
Table 3 shows the results of different methods on extractive datasets, where statistical (Knight and Marcu, 2000; Unno et al., 2006), ILP (Clarke and Lapata, 2008; Thadani and McKeown, 2013; Thadani, 2014) and dependency parsing (Cohn and Lapata, 2009) approaches as well as neural models such as RNNs (Filippova et al., 2015; Zhao et al., 2018; Kamigaito and Okumura, 2020) and Transformers (Park et al., 2021; Ghalandari et al., 2022) are reported.
@@
In specific, for the Ziff-Davis corpus, the highest relations F1 (56.55%) is obtained with a dependency tree that prunes unnecessary words from the input (Cohn and Lapata, 2009). The highest F-measure (78.58%) is achieved with a noisy channel model (Unno et al., 2006) that is extended with a maximum entropy model. Note that approaches beyond 2013 for the Ziff-Davis corpus are not included in Table 3 as recent methods typically employ larger datasets for evaluating their performance.
@@
For the Google sentence compression dataset, we can infer from Table 3 that the incorporation of syntactic and linguistic features boosts the performance of the different neural models (Kamigaito and Okumura, 2020; Park et al., 2021) compared to the baseline LSTM model (Filippova et al., 2015). The combination of dependency parsing techniques with a Transformer-base model (Park et al., 2021) yields the best results with more than 86 % in terms of F1 score. In addition, the introduction of RL (Zhao et al., 2018) also improves the performance of the baseline model (85.1% compared to 82% in terms of F1 score).
@@
BNC and Broadcast datasets have been extensively used for the evaluation of ILP methods. For the Broadcast dataset, an ILP method with Lagrangian relaxation (Thadani, 2014) shows additional gains against traditional ILP methods with global constraints (Clarke and Lapata, 2008) or commodity flow constraints (Thadani and McKeown, 2013). On the other hand, for the BNC dataset, an unsupervised Transformer-based method (Ghalandari et al., 2022) that employs RL achieves the highest F1 score of all the ILP methods. In addition, a higher compression (76% compared to 68.5%) is noticed, which might indicate reduced information loss during the generation of the compression.
@@
Table 4. Results of sentence compression methods on different abstractive datasets. We report ROUGE-1 (R-1), ROUGE-2 (R-2) and ROUGE-L (R-L) scores and compression rate (CR). All scores are reported in %. The first section of each dataset refers to supervised methods, while the second section, refers to unsupervised methods. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
Gigaword (Rush et al., 2015) 31.00 12.65 28.34 -
(Chopra et al., 2016) 33.78 15.97 31.15 -
(Zheng et al., 2020) 40.03 17.93 36.74 -
(Guan, 2021) 44.42 31.07 40.71 -
(Baziotis et al., 2019) 25.39 8.21 22.68 -
(Zhou and Rush, 2019) 26.48 10.05 24.41 -
(Malireddy et al., 2020) 29.80 7.52 26.10 -
(Ghalandari et al., 2022) 29.64 9.98 26.57 28.0
DUC 2004 (Rush et al., 2015) 28.18 8.49 23.81 -
(Chopra et al., 2016) 28.97 8.26 24.06 -
(Guan et al., 2021b) 30.03 10.71 26.05 -
(Zheng et al., 2020) 34.69 14.39 31.07 -
(Baziotis et al., 2019) 22.13 6.18 19.3 -
(Malireddy et al., 2020) 22.92 5.52 19.85 -
(Ghalandari et al., 2022) 25.27 7.82 22.14 35.0
(Schumann et al., 2020) 26.04 8.06 22.90 -
(Liu et al., 2022a) 26.71 7.68 23.06 -
(Hyun et al., 2022) 27.88 9.35 24.49 -
@@
Table 4 shows the results of different methods on abstractive sentence compression datasets, where different neural models are reported. Among all the supervised methods for the Gigaword dataset, the feed-forward network with a convolutional attention-based encoder (Rush et al., 2015) achieves a moderate score of 31% in terms of ROUGE-1. Even though this method outperforms all the unsupervised approaches (Baziotis et al., 2019; Zhou and Rush, 2019; Malireddy et al., 2020; Ghalandari et al., 2022), it fails to compete with the supervised RNN-based methods (Chopra et al., 2016; Zheng et al., 2020; Guan, 2021). More specifically, the attention-based RNN (Chopra et al., 2016) demonstrates a slight improvement with 33.78% ROUGE-1 which indicates that indeed RNNs are more powerful in this context, allowing for capturing longer dependencies and as a result generating more relevant compressions. An important boost in performance is noticed when LSTMs are paired with controllable attributes such as guiding entities (Zheng et al., 2020), resulting in an increase of approximately 7% points in terms of the ROUGE-1 score. The same conclusions are drawn with the incorporation of knowledge graphs (Guan, 2021) with a significant increase in performance of 44.42% and 31.07% in terms of ROUGE-1 and ROUGE-2 scores, respectively. Regarding the unsupervised approaches for the Gigaword dataset, the LSTM-based method with the compressor and reconstructor (Baziotis et al., 2019) achieves a relatively low ROUGE-1 score of 25.39%, while slight improvements are noticed when ELMo embeddings are added in a similar LSTM architecture (Zhou and Rush, 2019). The approach of (Malireddy et al., 2020), which optimizes the LSTM architecture to not only reconstruct but also rephrase words within the text, demonstrates additional gains with a noticeable improvement of 29.80% ROUGE-1 score. In addition, Transformers with RL (Ghalandari et al., 2022) achieve a very similar ROUGE-1 score (29.64%) but a slightly higher ROUGE-L score (26.57%). This might indicate that while Transformers are powerful in many NLP tasks, RNN architectures are still useful in this context.
@@
Similar trends are observed on the DUC 2004 dataset, where the same supervised and unsupervised techniques are reported. The supervised approaches, such as the feed-forward neural network (Rush et al., 2015), attention-based RNN (Chopra et al., 2016) and the LSTM with knowledge graphs (Guan, 2021) demonstrate comparable performance with approximately $30\%$ ROUGE-1 score, while the LSTM with guiding entities outperforms all the previous approaches with more than 34% ROUGE-1 score. Among the unsupervised approaches, we notice the same patterns with the Gigaword dataset, while there are slight improvements specifically with the incorporation of discrete optimization techniques (Schumann et al., 2020; Liu et al., 2022a) and the use of RL with Transformers (Hyun et al., 2022). Since DUC 2004 is typically used only for evaluation, the ROUGE scores for both the unsupervised and supervised methods remain relatively low compared to the Gigaword dataset. Similar to extractive models in Table 3, we notice again that the incorporation of additional features and techniques into sequence-to-sequence models can improve the overall performance of compression models.
@@
The overall comparative evaluation demonstrates that methods that employ Transformer-based architectures (Hyun et al., 2022; Ghalandari et al., 2022) seems to outperform the majority of existing methods for both supervised and unsupervised models. On the other hand, in some cases, LSTM architectures can compete with Transformers, achieving similar ROUGE scores in an unsupervised setting. Additional gains are observed when Transformers are combined with other techniques such as RL or controllable attributes i.e., length and entities. These results also indicate the potential benefits of different attribute controls in generating more accurate and relevant compressions. Finally, we notice that methods on extractive datasets achieve overall higher ROUGE scores than methods on abstractive datasets. Extractive methods have a significant advantage over ROUGE scores since they are based on deletion-only operations, in contrast to abstractive methods, which are based on rewriting or paraphrasing operations.
3.5. Other Trends

### v2

@@ 3.2.1 Extractive Datasets
The first dataset for extractive sentence compression is extracted from the Ziff-Davis corpus Knight2000 . Ziff-Davis has a relatively small corpus of newspaper articles announcing computer products. A small portion of the corpus has about 1,000 sentence-compression pairs Knight2002 . This dataset has been mostly used for training and evaluating early works on sentence compression. In addition,  vandeghinste2004using build a parallel corpus for subtitle generation that contains transcriptions from several broadcasts. The British National Corpus (BNC) and the English Broadcast News Corpus are two other small human-annotated extractive sentence compression datasets that are typically used for evaluation purposes Clarke2008GlobalApproach . The first one comprises 82 newspaper articles (1,433 sentences) from the BNC and the American News Text corpus, while the second one comprises 50 stories (1,370 sentences) from the HUB-4 1996 English Broadcast News corpus.
@@
Summarization datasets have also been used for the evaluation of sentence compression. For example, a small extractive dataset of 280 examples is constructed in Ghalandari2022EfficientLearning based on Newsroom. Newsroom Grusky2018Newsroom:Strategies is a large dataset for training and evaluating summarization systems. It contains 1.3 million articles and summaries written by authors and editors in the newsrooms of 38 major publications. The summaries are obtained from search and social metadata between 1998 and 2017 and use a variety of summarization strategies combining extraction and abstraction. Ghalandari2022EfficientLearning extract the first three sentences from each article with a length between 15 and 60 tokens. They consider cases where the tokenized summary was contained in a tokenized sentence. They further filter these cases based on grammaticality and informativeness.
@@
Google sentence compression Filippova2013OvercomingCompression is the largest dataset for extractive sentence compression. It is automatically constructed, consisting of 250K first sentences of a news article along with their extractive compression derived from information in the corresponding article’s headline [^1] https://github.com/google-research-datasets/sentence-compression.
3.2.2 Abstractive Datasets
@@
Abstractive datasets contain sentence-compression pairs, where the compression is not necessarily a direct sub-sequence of the original long sentence. While the majority of related literature has focused on the news domain, there also exists work in other domains, such as subtitles or Wikipedia documents. For example, yamangil2008mining exploit Wikipedia revision history to enrich compression data with more than 380,000 compression pairs.
@@
A small abstractive dataset is constructed in Rush2015ASummarization based on the data of Task 1 and 2 of the 2004 Document Understanding Conferences (DUC) [^2] https://duc.nist.gov/duc2004/. It contains 500 news articles from the New York Times and Associated Press Wire services, each paired with four human summaries. A repository containing these data is available [^3] https://github.com/UsmanNiazi/DUC-2004-Dataset. For the purpose of sentence summarization, the first sentence of each article is used, while summaries are clipped at 75 characters. This leads to summaries with an average of 10 words.
@@
A larger abstractive text compression dataset toutanova-etal-2016-dataset is introduced from Microsoft Research (MSR). It consists of 26,000 long sentences accompanied by their respective compressions. Each long sentence is paired with multiple target compressions (1 to 5) rated by quality judgments regarding sentence coherence and meaning preservation. Besides the newswire domain, the MSR dataset also includes journals as well as business and technical reports. One of the largest datasets for abstractive sentence compression is Gigaword Napoles2012 . Gigaword is used in Rush2015ASummarization to create a dataset by pairing the first sentence of each article with its headline. Gigaword is a large-scale dataset that is used for summarization and headline generation. It contains approximately 9.5 million articles from different news sources. After rule-based filtering of some spurious pairs, about 4 million examples remained.
@@
Abstractive datasets for sentence compression are either used only for evaluation (DUC-2004) while larger datasets such as Gigaword Napoles2012 and MSR toutanova-etal-2016-dataset are also used for training and evaluation.
3.3 Measures
@@
3.3.1 Compression Rate
One of the most widely used metric for evaluating sentence compression methods is Compression Rate Knight2000 ; McDonald2006 ; Galley2007 ; Galanis2010 ; Filippova2015 ; zhou2019 ; Zhao2018 ; Ghalandari2022EfficientLearning . It is defined as the number of words in the compressed text divided by the number of tokens in the uncompressed text as follows:
[…]
@@ 3.3.2 ROUGE score
The family of ROUGE metrics Lin2004Rouge is one of the most popular measures for text summarization. Since text summarization and sentence compression share many similarities in terms of shortening the initial text into a more concise form, ROUGE is also widely used for sentence compression. ROUGE metrics are based on the n-gram overlaps or the longest common subsequence (LCS) between the target and the prediction and can be used to evaluate the quality of the generated text. ROUGE consists of several metrics such as ROUGE-N (ROUGE-1, ROUGE-2, etc.), ROUGE-L, and ROUGE-S. In sentence compression, ROUGE metrics are computed between the predicted and ground truth compressions. More specifically, ROUGE-1 and ROUGE-2 are used for measuring the overlap of unigrams and bigrams between the predicted and ground truth compression, respectively. ROUGE-L is used for measuring the LCS between the generated and the target compression.
@@
Even though ROUGE is an established evaluation metric for sentence compression, Zhang2021Fact highlight some of its limitations. For example, high ROUGE score might be misleading producing factual errors. li2018ensure reveal that existing encoder-decoder models zhou2017selective that achieve significant performance according to established metrics like ROUGE score might not generate faithful compressions. Another important factor in this context is the informativeness of the generated compressions. zheng2020controllable work towards this direction by proposing an entity-controllable compression model that ensures the presence of the most relevant entities of the input sentence in the compressions. This method works by first extracting all the entities from the input sentence and ranking the most important ones with an entity selector module. Finally, instead of conventional left-to-right decoding, the model decodes the text around the selected entities, ensuring that the resulting compression is both concise and informative.
3.3.3 Other measures
@@
Other measures include simple token-overlapping Ghalandari2022EfficientLearning , grammatical relations clarke2006models ; Filippova2013OvercomingCompression or even human evaluation Filippova2013OvercomingCompression . In specific, clarke2006models demonstrate that measuring the grammatical relations between the input and the target is a measure that correlates well with human judgments. The F1 score between the functional structures of the sentence and its compression, obtained via Lexical-Functional Grammar parsing, was proposed in Riezler2003StatisticalGrammar and used in Filippova2013OvercomingCompression . In addition, simple F1 score between the tokens of the predicted and ground truth compression are computed in Ghalandari2022EfficientLearning . In Filippova2013OvercomingCompression , humans were shown pairs of sentences and their compression and were asked to rate in a scale from 1 to 5 the readability (i.e., grammaticallity, fluency) and the informativeness (i.e., importance, representatives) of the compression. Three ratings were collected for every compression.
3.4 Comparative Evaluation
@@
Table 3: Results of sentence compression methods on different extractive datasets. We report ROUGE-1 (R-1), ROUGE-2 (R-2) and ROUGE-L (R-L) scores, F-measure, relations F-measure (Relations F1) and compression rate (CR). All scores are reported in %. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
Ziff-Davis Knight2000 - - - 68.80 - 70.19
unno2006trimming - - - 78.58 - 58.14
Clarke2008GlobalApproach - - - - 46.77 48.67
Cohn2009 - - - - 56.55 67.45
Google Filippova2015 - - - 82.0 - 38.0
Zhao2018 - - - 85.1 - 39.0
kamigaito2020syntactically 79.3 71.4 79.1 85.5 - -
park2021sentence - - - 86.2 - 40.9
BNC Clarke2008GlobalApproach - - - - 50.1 68.5
thadani2013-sentence - - - 71.91 - -
thadani2014approximation - - - 74.00 - -
Ghalandari2022EfficientLearning 79.49 62.32 78.63 76.50 - 76.0
Broadcast Clarke2008GlobalApproach - - - 40.8 63.7
Ghalandari2022EfficientLearning 82.22 66.01 81.78 78.7 - 78.0
thadani2013-sentence - - - 77.82 - -
thadani2014approximation - - - 80.13 - -
@@
Table 3 shows the results of different methods on extractive datasets, where statistical Knight2000 ; unno2006trimming , ILP Clarke2008GlobalApproach ; thadani2013-sentence ; thadani2014approximation and dependency parsing Cohn2009 approaches as well as neural models such as RNNs Filippova2015 ; Zhao2018 ; kamigaito2020syntactically and Transformers park2021sentence ; Ghalandari2022EfficientLearning are reported.
@@
In specific, for the Ziff-Davis corpus, the highest relations F1 (56.55%) is obtained with a dependency tree that prunes unnecessary words from the input Cohn2009 . The highest F-measure (78.58%) is achieved with a noisy channel model unno2006trimming that is extended with a maximum entropy model. Note that approaches beyond 2013 for the Ziff-Davis corpus are not included in Table 3 as recent methods typically employ larger datasets for evaluating their performance.
@@
For the Google sentence compression dataset, we can infer from Table 3 that the incorporation of syntactic and linguistic features boosts the performance of the different neural models kamigaito2020syntactically ; park2021sentence compared to the baseline LSTM model Filippova2015 . The combination of dependency parsing techniques with a Transformer-base model park2021sentence yields the best results with more than 86 % in terms of F1 score. In addition, the introduction of RL Zhao2018 also improves the performance of the baseline model (85.1% compared to 82% in terms of F1 score).
@@
BNC and Broadcast datasets have been extensively used for the evaluation of ILP methods. For the Broadcast dataset, an ILP method with Lagrangian relaxation thadani2014approximation shows additional gains against traditional ILP methods with global constraints Clarke2008GlobalApproach or commodity flow constraints thadani2013-sentence . On the other hand, for the BNC dataset, an unsupervised Transformer-based method Ghalandari2022EfficientLearning that employs RL achieves the highest F1 score of all the ILP methods. In addition, a higher compression (76% compared to 68.5%) is noticed, which might indicate reduced information loss during the generation of the compression.
@@
Table 4: Results of sentence compression methods on different abstractive datasets. We report ROUGE-1 (R-1), ROUGE-2 (R-2) and ROUGE-L (R-L) scores and compression rate (CR). All scores are reported in %. The first section of each dataset refers to supervised methods, while the second section, refers to unsupervised methods. Results are obtained from the corresponding papers. Dashes (-) indicate that the score is not provided.
@@
Gigaword Rush2015ASummarization 31.00 12.65 28.34 -
Chopra2016 33.78 15.97 31.15 -
zheng2020controllable 40.03 17.93 36.74 -
guan2021knowledge 44.42 31.07 40.71 -
Baziotis2019 25.39 8.21 22.68 -
zhou2019 26.48 10.05 24.41 -
Malireddy2020 29.80 7.52 26.10 -
Ghalandari2022EfficientLearning 29.64 9.98 26.57 28.0
DUC 2004 Rush2015ASummarization 28.18 8.49 23.81 -
Chopra2016 28.97 8.26 24.06 -
guan2021integrating 30.03 10.71 26.05 -
zheng2020controllable 34.69 14.39 31.07 -
Baziotis2019 22.13 6.18 19.3 -
Malireddy2020 22.92 5.52 19.85 -
Ghalandari2022EfficientLearning 25.27 7.82 22.14 35.0
Schumann2020DiscreteExtraction 26.04 8.06 22.90 -
liu2022learning 26.71 7.68 23.06 -
hyun2022generating 27.88 9.35 24.49 -
@@
Table 4 shows the results of different methods on abstractive sentence compression datasets, where different neural models are reported. Among all the supervised methods for the Gigaword dataset, the feed-forward network with a convolutional attention-based encoder Rush2015ASummarization achieves a moderate score of 31% in terms of ROUGE-1. Even though this method outperforms all the unsupervised approaches Baziotis2019 ; zhou2019 ; Malireddy2020 ; Ghalandari2022EfficientLearning , it fails to compete with the supervised RNN-based methods Chopra2016 ; zheng2020controllable ; guan2021knowledge . More specifically, the attention-based RNN Chopra2016 demonstrates a slight improvement with 33.78% ROUGE-1 which indicates that indeed RNNs are more powerful in this context, allowing for capturing longer dependencies and as a result generating more relevant compressions. An important boost in performance is noticed when LSTMs are paired with controllable attributes such as guiding entities zheng2020controllable , resulting in an increase of approximately 7% points in terms of the ROUGE-1 score. The same conclusions are drawn with the incorporation of knowledge graphs guan2021knowledge with a significant increase in performance of 44.42% and 31.07% in terms of ROUGE-1 and ROUGE-2 scores, respectively. Regarding the unsupervised approaches for the Gigaword dataset, the LSTM-based method with the compressor and reconstructor Baziotis2019 achieves a relatively low ROUGE-1 score of 25.39%, while slight improvements are noticed when ELMo embeddings are added in a similar LSTM architecture zhou2019 . The approach of Malireddy2020 , which optimizes the LSTM architecture to not only reconstruct but also rephrase words within the text, demonstrates additional gains with a noticeable improvement of 29.80% ROUGE-1 score. In addition, Transformers with RL Ghalandari2022EfficientLearning achieve a very similar ROUGE-1 score (29.64%) but a slightly higher ROUGE-L score (26.57%). This might indicate that while Transformers are powerful in many NLP tasks, RNN architectures are still useful in this context.
@@
Similar trends are observed on the DUC 2004 dataset, where the same supervised and unsupervised techniques are reported. The supervised approaches, such as the feed-forward neural network Rush2015ASummarization , attention-based RNN Chopra2016 and the LSTM with knowledge graphs guan2021knowledge demonstrate comparable performance with approximately $30\%$ ROUGE-1 score, while the LSTM with guiding entities outperforms all the previous approaches with more than 34% ROUGE-1 score. Among the unsupervised approaches, we notice the same patterns with the Gigaword dataset, while there are slight improvements specifically with the incorporation of discrete optimization techniques Schumann2020DiscreteExtraction ; liu2022learning and the use of RL with Transformers hyun2022generating . Since DUC 2004 is typically used only for evaluation, the ROUGE scores for both the unsupervised and supervised methods remain relatively low compared to the Gigaword dataset. Similar to extractive models in Table 3, we notice again that the incorporation of additional features and techniques into sequence-to-sequence models can improve the overall performance of compression models.
@@
The overall comparative evaluation demonstrates that methods that employ Transformer-based architectures hyun2022generating ; Ghalandari2022EfficientLearning seems to outperform the majority of existing methods for both supervised and unsupervised models. On the other hand, in some cases, LSTM architectures can compete with Transformers, achieving similar ROUGE scores in an unsupervised setting. Additional gains are observed when Transformers are combined with other techniques such as RL or controllable attributes i.e., length and entities. These results also indicate the potential benefits of different attribute controls in generating more accurate and relevant compressions. Finally, we notice that methods on extractive datasets achieve overall higher ROUGE scores than methods on abstractive datasets. Extractive methods have a significant advantage over ROUGE scores since they are based on deletion-only operations, in contrast to abstractive methods, which are based on rewriting or paraphrasing operations.
3.5 Other Trends